OpenStudy (anonymous):

6 6/7 = n + 5 3/7 ? a) 12 2/7 b) 6 6/7 c) 5 6/7 d) 1 3/7 I will give a medal

OpenStudy (matlee):

Hes hot

OpenStudy (matlee):

Tbing trade phone numbers wit me

OpenStudy (matlee):

ur so cute how do you do ur hair

OpenStudy (anonymous):

I was actually born with this color hair

OpenStudy (anonymous):

Weird right

OpenStudy (matlee):

thats cool i dyed mine ur is much more pretty though

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

so what doe the question ask for? to solve for N?

OpenStudy (anonymous):

WILL FAN AND GIVE METAL IF ANSWER IS RIGHT!!!Match each pair of points with the coordinate that will make a right triangle with its hypotenuse along the graph shown. http://static.k12.com/calms_media/media/1503500_1504000/1503535/1/cf999c9f1c44ae152ac410ef8d736e7374780828/MS_IMC-140523-130904.jpg Column A Column B 1. (2, 3) and (4, 6) 2. (4, 6) and (8, 12) 3. (2, 3) and (8, 12) 4. (2, 3) and (6, 9) A. (8, 3) B. (4, 3) C. (8, 6) D. (6, 6) E. (6, 3)

OpenStudy (greencat):

\[6 \frac{ 6 }{ 7 }=n + \frac{3}{7}\] \[n=5\frac{3}{7}-6 \frac{6}{7}\]

OpenStudy (anonymous):

To Figure out what N is @Dfoeltz

OpenStudy (anonymous):

looks like green cats got ya covered! haha

OpenStudy (anonymous):

CAN SOMEONE HELP ME

OpenStudy (anonymous):

PLZ

OpenStudy (anonymous):

UHH WHAT IS GOING ON

OpenStudy (greencat):

Make the fractions improper. So: 5 3/7= 6 6/7=

OpenStudy (anonymous):

heywassup your on the wrong thing post in the entirety of of mathematics help not a question

OpenStudy (anonymous):

WILL FAN AND GIVE METAL IF ANSWER IS RIGHT!!!Match each pair of points with the coordinate that will make a right triangle with its hypotenuse along the graph shown. http://static.k12.com/calms_media/media/1503500_1504000/1503535/1/cf999c9f1c44ae152ac410ef8d736e7374780828/MS_IMC-140523-130904.jpg Column A 1. (2, 3) and (4, 6) 2. (4, 6) and (8, 12) 3. (2, 3) and (8, 12) 4. (2, 3) and (6, 9) Column B A. (8, 3) B. (4, 3) C. (8, 6) D. (6, 6) E. (6, 3)

OpenStudy (greencat):

Just make them improper franctions and subtract.

OpenStudy (anonymous):

MATLEE SHUT UP NOBODY WANTS UR NUMBER

OpenStudy (anonymous):

OH OK

OpenStudy (anonymous):

AND THAT MY ANSWER FOR ALL OF THEM

OpenStudy (anonymous):

WHAT IS THE QUESTION

OpenStudy (anonymous):

CAN U DO THE FIRST ONE AS AN EXAMPLE PLZ

OpenStudy (anonymous):

WHO IS@MATLEE

OpenStudy (anonymous):

D,A,C,B,C,A,D,C,B

OpenStudy (greencat):

tbing u can solve the rest. it should be easy. bye